# The Butterfly Effect of Altering Prompts: How Small Changes and Jailbreaks Affect Large Language Model Performance

### v2

@@ 4.2 Do prompt variations affect accuracy?
Table 1: Accuracy of each prompt variation on each task. Red percentages indicate that the accuracy dropped from there baseline (Python List Format) while green percentages indicate the accuracy increased.
Yes! Table 1 shows the accuracies of each prompt variation across our 11 tasks. We find that there isn’t a one-size-fits-all formatting specification or perturbation suitable for every task, although we generally observe success using the Python List, No Specified Format, or JSON specification. No Specified Format leads to the overall most accurate results, beating the next best variation by a whole percentage point.
Formatting in YAML, XML, or CSV leads to a 3-6% loss in accuracy in comparison to the Python List specification, with CSV displaying the lowest performance across all output formats. Surprisingly, CSV achieves the highest accuracy among all variations for the IMDBSentiment task, albeit by only a marginal percentage point. This emphasizes the absence of a definitive “best” output format for usage.
When it comes to influencing the model by specifying a tip versus specifying we will not tip, we noticed minimal performance changes. Tipping and not tipping resulted in the same overall accuracy. Contrary to expectations, even tipping an extravagant amount, specifically \$1000, did not lead to a substantial difference in accuracy.
Furthermore, our experimentation revealed a significant performance drop when using certain jailbreaks. AIM and Dev Mode v2 unsurprisingly exhibit very low accuracy, primarily due to a majority of their responses being invalid. Similarly, Evil Confidant, with its prompt guiding it toward “unhinged” responses, also yields low accuracy. Surprisingly, the Refusal Suppression alone results in a loss of over 10% accuracy (compared to Python List), highlighting the inherent instability even in seemingly innocuous jailbreaks. This underscores the unpredictability associated with jailbreak usage.
4.3 How similar are the predictions from each prompt variation?
Figure 3: MDS representation of all variation’s predictions.
First, we observe an interesting relationship between Python List specification and the No Specified Format. These two vectors are placed close together in the MDS representation. We note again that these two formats also achieved the highest overall accuracy. Adjacent to these points were simple perturbations, which were formatted as Python lists, such as initial greetings or the addition of a space. This clustering around the Python List variation may be attributed to these prompts having only a few token differences while preserving the overall semantics.
Contrary to expectations, all tipping variations clustered together, including the Won’t Tip variation. Surprisingly, increasing the tip amount exhibited a linear relationship with distances from the Won’t Tip variation.
End with “Thank you” and Rephrase as Statement stood out as outliers, situated far from the main clusters. It is surprising that simply thanking the model can lead to such a considerable difference while adding a greeting or space token leads to a minimal change. The substantial impact of rephrasing was expected, given the increased token changes compared to other prompts.
Lastly, the jailbreak variations displayed a wider spread, and were expectedly away from the most accurate changes. Dev Mode v2 and AIM predominantly produced invalid responses, aligning with their broader distribution. Surprisingly, Refusal Suppression fell on the outskirts of the primary cluster, possibly due to the extensive token addition through the jailbreak. The Evil Confidant variation, despite requiring fewer tokens, notably diverged from the cluster due to its distinct “unhinged” responses.
Category Correlation
All -0.2311 (p = 0.00)
Output Formats -0.0720 (p = 0.02)
Perturbations 0.1209 (p = 0.00)
Tipping 0.1241 (p = 0.00)
Jailbreaks -0.3779 (p = 0.00)
4.4 Do variations correlate to annotator disagreement?
Not really! Leveraging the Jigsaw Toxicity task, which we specifically sampled only to include samples with 100 or more annotations, we hypothesized that more confusing samples would lead to more annotator disagreement and more variation in ChatGPT’s predictions. To aid our analysis, we calculate the entropy of annotator predictions and the entropy of our predictions per sample.
Table 2 lists the Pearson correlations between the Jigsaw Toxicity predictions, across each category of prompt variations. We identify some weak correlations with annotator disagreement. However, the strongest correlations are negative, meaning that the most confusing instances (i.e., lowest entropy) and the most likely to change. This indicates that the confusion of the instance provides some explanatory power for why the prediction changes, but there are other factors at play.
5 Conclusion
In this paper, we investigate how simple and commonly-used prompt variations can affect an LLM’s predictions. We demonstrate that even minor prompt variations can change a considerable proportion of predictions. That said, despite some fraction of labels changing, most perturbations yield similar accuracy. We find that jailbreaks lead to considerable performance losses. The AIM and Dev Mode v2 jailbreaks led to refusal rates around 90%. Additionally, while both Evil Confidant and Refusal Suppression had a refusal rate of less than 3%, their inclusion led to a loss of over 10 percentage points compared to our baseline. Finally, we observe a performance hit of, on average, 5% when using specific output format specifications, namely CSV, XML, and ChatGPT’s JSON Checkbox.
Next we analyze the patterns of these changes. First, we embed the prompt variations based on their subsequent responses using MDS, and find that perturbation outputs tend to more closely resemble our baseline than formatting changes, and that both have higher fidelity than jailbreaks. Next, we study the correlation between annotator disagreement and an instance’s propensity to change. We find a slight correlation between annotator disagreement and the likelihood to change.
The directions for future work is abundant. A major next step would be to generate LLMs that are resilient to these changes, offering consistent answers across formatting changes, perturbations, and jailbreaks. Towards that goal, future work includes seeking a firmer understanding of why responses change under minor changes to the prompt, and better anticipating an LLMs change in its response to a particular instance.
Acknowledgements
This project was sponsored by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR00112290021. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA).
Appendix A Full Prompts
A.1 Tasks
Each task and a corresponding example prompt is shown in Table 3.
Task Prompt
Table 3: Examples of each task’s prompt.
A.2 Variations
Table 4: Examples of each variation’s prompt.
Variation Example
Appendix B Additional Experimental Results
In our paper, we discussed how many overall predictions change when a prompt variation is used. For completeness, Table 5 and  6 include these results on a per task level.
@@ Appendix B Additional Experimental Results
Table 5: Number of labels changed compared to No Specified Format per task for each output format.
@@
Table 6: Number of labels changed compared to Python List per task for each variation.
@@ Appendix C No Specified Format Analysis
Figure 4: Number of predictions that change (out of 11,000) compared to the No Specified Format. Red bars correspond to the number of invalid responses provided by the model.
@@
Table 7: Accuracy of each prompt variation on each task when using no specified output format on each variation. Red percentages indicate that the accuracy dropped from there baseline (Python List Format) while green percentages indicate the accuracy increased.

### v3

@@ Appendix B Extended Results
Table 5: ChatGPT’s Accuracy of each prompt variation on each task. Red percentages indicate that the accuracy dropped from there baseline (Python List Format) while green percentages indicate the accuracy increased.
BoolQ
CoLA
ColBERT
CoPA
GLUE Diagnostic
IMDBSentiment
iSarcasm
Jigsaw Toxicity
MathQA
RACE
TweetStance
Overall
Python List Format 67% 43% 21% 33% 35% 72% 23% 67% 20% 48% 29% 42%
JSON Format 70% 66% 28% 44% 29% 84% 27% 64% 14% 47% 33% 46%
XML Format 53% 57% 25% 53% 40% 72% 21% 62% 14% 52% 31% 44%
CSV Format 67% 39% 19% 39% 38% 76% 25% 60% 20% 48% 33% 42%
YAML Format 61% 64% 22% 40% 26% 69% 28% 65% 16% 49% 37% 44%
No Specified Format 48% 55% 18% 42% 34% 79% 37% 68% 8% 49% 26% 42%
Start with "Hello." 42% 64% 33% 26% 40% 75% 30% 67% 11% 47% 38% 43%
Start with "Hello!" 45% 66% 34% 30% 36% 78% 27% 68% 13% 47% 38% 44%
Start with "Howdy!" 41% 58% 26% 28% 33% 73% 24% 64% 12% 47% 32% 40%
End with "Thank you" 67% 45% 22% 30% 35% 77% 25% 68% 23% 48% 35% 43%
Rephrase as Statement 61% 52% 28% 57% 37% 78% 51% 70% 24% 50% 35% 49%
Won’t Tip 63% 32% 15% 18% 30% 62% 20% 57% 21% 43% 28% 35%
Tip \$1 69% 68% 38% 58% 41% 85% 33% 64% 25% 50% 42% 52%
Tip \$10 69% 69% 42% 58% 42% 85% 33% 64% 25% 48% 43% 53%
Tip \$100 69% 67% 37% 51% 42% 83% 33% 62% 24% 49% 40% 51%
Tip \$1000 69% 59% 33% 40% 43% 80% 32% 59% 24% 48% 38% 48%
AIM 18% 29% 6% 4% 4% 47% 5% 26% 1% 29% 44% 19%
Evil Confidant 24% 62% 11% 5% 22% 48% 23% 51% 10% 30% 32% 29%
Refusal Suppression 62% 49% 19% 54% 44% 81% 14% 49% 23% 47% 27% 43%
Dev Mode v2 31% 37% 24% 9% 23% 41% 21% 34% 14% 13% 43% 26%
Aggregate Styles 72% 66% 22% 49% 36% 81% 26% 74% 21% 51% 28% 48%
Aggregate Perturbations 65% 59% 24% 35% 37% 78% 28% 71% 21% 49% 31% 45%
Aggregate Jailbreaks 38% 59% 17% 13% 24% 74% 15% 38% 16% 35% 38% 33%
Aggregate Tipping 69% 68% 35% 52% 43% 84% 33% 64% 28% 49% 38% 51%
Table 6: Llama 7B’s accuracy of each prompt variation on each task. Red percentages indicate that the accuracy dropped from there baseline (Python List Format) while green percentages indicate the accuracy increased.
BoolQ
CoLA
ColBERT
CoPA
GLUE Diagnostic
IMDBSentiment
iSarcasm
Jigsaw Toxicity
MathQA
RACE
TweetStance
Overall
Python List Format 74% 64% 42% 73% 46% 91% 48% 66% 43% 54% 35% 58%
JSON Format 64% 60% 43% 77% 47% 91% 47% 68% 35% 56% 31% 56%
XML Format 76% 68% 35% 65% 47% 84% 43% 56% 43% 55% 30% 55%
CSV Format 72% 71% 42% 76% 47% 91% 48% 51% 45% 55% 33% 57%
YAML Format 72% 67% 47% 77% 47% 77% 50% 63% 42% 55% 34% 57%
No Specified Format 73% 58% 48% 65% 46% 90% 47% 41% 43% 50% 31% 54%
Start with "Hello." 73% 58% 39% 69% 44% 90% 52% 52% 38% 55% 32% 55%
Start with "Hello!" 72% 62% 42% 70% 45% 91% 52% 55% 38% 56% 34% 56%
Start with "Howdy!" 72% 57% 42% 65% 46% 90% 54% 51% 39% 54% 31% 55%
End with "Thank you" 74% 64% 37% 70% 45% 92% 49% 59% 40% 55% 35% 57%
Rephrase as Statement 68% 60% 43% 75% 44% 91% 35% 53% 41% 57% 31% 54%
Won’t Tip 72% 63% 29% 66% 45% 92% 46% 64% 37% 55% 37% 55%
Tip \$1 73% 69% 36% 70% 47% 92% 55% 68% 36% 56% 35% 58%
Tip \$10 71% 66% 31% 69% 46% 93% 54% 65% 36% 55% 32% 56%
Tip \$100 71% 63% 26% 67% 45% 93% 53% 58% 37% 55% 28% 54%
Tip \$1000 69% 61% 20% 66% 44% 91% 51% 53% 38% 54% 24% 52%
AIM 49% 36% 19% 8% 23% 83% 22% 35% 2% 42% 12% 30%
Evil Confidant 32% 27% 7% 3% 6% 53% 13% 45% 2% 22% 17% 21%
Refusal Suppression 65% 70% 49% 69% 38% 71% 46% 64% 39% 54% 38% 55%
Dev Mode v2 41% 66% 54% 17% 38% 83% 34% 74% 26% 32% 44% 46%
Aggregate Styles 75% 64% 46% 76% 47% 92% 46% 68% 48% 55% 32% 59%
Aggregate Perturbations 74% 61% 41% 72% 46% 92% 50% 58% 47% 56% 31% 57%
Aggregate Jailbreaks 47% 54% 23% 5% 23% 87% 24% 63% 7% 45% 20% 36%
Aggregate Tipping 71% 64% 30% 68% 46% 93% 53% 65% 39% 55% 30% 56%
Table 7: Llama 13B’s accuracy of each prompt variation on each task. Red percentages indicate that the accuracy dropped from there baseline (Python List Format) while green percentages indicate the accuracy increased.
BoolQ
CoLA
ColBERT
CoPA
GLUE Diagnostic
IMDBSentiment
iSarcasm
Jigsaw Toxicity
MathQA
RACE
TweetStance
Overall
Python List Format 81% 68% 42% 83% 37% 92% 71% 77% 37% 72% 54% 65%
JSON Format 86% 75% 54% 79% 40% 93% 77% 86% 34% 64% 69% 69%
XML Format 83% 64% 33% 78% 31% 86% 36% 66% 23% 65% 52% 56%
CSV Format 75% 78% 52% 77% 23% 82% 73% 72% 45% 68% 57% 64%
YAML Format 83% 70% 45% 78% 35% 61% 69% 72% 48% 73% 42% 61%
No Specified Format 86% 62% 58% 65% 38% 90% 61% 85% 46% 60% 66% 65%
Start with "Hello." 82% 69% 41% 82% 37% 92% 69% 72% 37% 73% 40% 63%
Start with "Hello!" 82% 70% 42% 83% 37% 92% 69% 74% 36% 72% 48% 64%
Start with "Howdy!" 79% 72% 40% 81% 35% 91% 68% 74% 37% 72% 40% 63%
End with "Thank you" 82% 60% 59% 83% 36% 90% 62% 77% 37% 72% 52% 64%
Rephrase as Statement 73% 72% 43% 81% 40% 93% 70% 78% 38% 72% 47% 64%
Won’t Tip 82% 61% 36% 80% 36% 93% 69% 76% 35% 73% 51% 63%
Tip \$1 78% 63% 42% 80% 37% 92% 69% 74% 35% 72% 41% 62%
Tip \$10 78% 62% 40% 76% 37% 92% 68% 75% 34% 72% 37% 61%
Tip \$100 77% 60% 35% 72% 34% 91% 66% 74% 35% 72% 33% 59%
Tip \$1000 77% 56% 30% 65% 32% 91% 64% 73% 36% 71% 29% 57%
AIM 78% 47% 51% 34% 40% 88% 40% 59% 35% 61% 71% 55%
Evil Confidant 24% 12% 4% 7% 6% 56% 4% 34% 6% 32% 12% 18%
Refusal Suppression 62% 71% 44% 74% 38% 67% 54% 65% 36% 61% 49% 57%
Dev Mode v2 53% 31% 41% 55% 21% 87% 32% 46% 25% 56% 48% 45%
Aggregate Styles 85% 78% 52% 81% 37% 93% 73% 84% 50% 69% 56% 69%
Aggregate Perturbations 81% 69% 42% 84% 37% 93% 70% 80% 37% 73% 50% 65%
Aggregate Jailbreaks 70% 52% 42% 60% 34% 89% 41% 59% 32% 65% 59% 55%
Aggregate Tipping 78% 61% 37% 76% 36% 92% 68% 76% 37% 73% 35% 61%
Table 8: Llama 70B’s accuracy of each prompt variation on each task. Red percentages indicate that the accuracy dropped from there baseline (Python List Format) while green percentages indicate the accuracy increased.
@@
Table 9: ChatGPT’s number of labels changed compared to No Specified Format per task for each output format.
CoLA
CoPA
ColBERT
GLUE Diagnostic
iSarcasm
IMDBSentiment
Jigsaw Toxicity
BoolQ
MathQA
RACE
TweetStance
Python List Format 548 529 695 436 661 242 387 452 843 233 577
JSON Format 360 472 674 635 621 171 379 498 787 261 590
XML Format 480 463 690 412 603 285 397 466 683 280 636
CSV Format 600 474 700 415 636 233 414 556 762 260 590
YAML Format 346 521 696 653 633 279 379 722 750 236 580
Table 10: Llama 7Bs number of labels changed compared to No Specified Format per dataset for each variation.
CoLA
CoPA
ColBERT
GLUE Diagnostic
iSarcasm
IMDBSentiment
TweetStance
Jigsaw Toxicity
BoolQ
MathQA
RACE
Python List Format 111 255 320 188 185 79 521 419 238 556 215
JSON Format 50 253 312 167 217 90 539 500 383 622 207
XML Format 329 341 613 194 249 144 575 579 277 517 224
CSV Format 640 235 336 185 263 72 525 569 242 524 198
YAML Format 593 231 301 152 308 247 516 513 260 558 211
Table 11: Llama 13Bs number of labels changed compared to No Specified Format per dataset for each variation.
CoLA
ColBERT
GLUE Diagnostic
iSarcasm
IMDBSentiment
Jigsaw Toxicity
MathQA
CoPA
TweetStance
BoolQ
RACE
Python List Format 128 463 455 531 64 263 646 305 405 89 262
JSON Format 352 277 267 378 54 181 647 319 303 121 292
XML Format 286 632 463 541 123 372 726 318 415 117 256
CSV Format 386 328 675 551 191 294 464 348 380 174 244
YAML Format 557 494 408 529 391 358 473 341 557 162 297
Table 12: Llama 70Bs number of labels changed compared to No Specified Format per dataset for each variation.
@@
Table 13: ChatGPT number of labels changed compared to Python List per task for each variation.
CoLA
CoPA
ColBERT
GLUE Diagnostic
iSarcasm
IMDBSentiment
Jigsaw Toxicity
BoolQ
MathQA
RACE
TweetStance
Start with "Hello." 504 428 439 383 497 259 294 504 808 175 454
Start with "Hello!" 517 443 464 397 515 233 281 439 800 160 481
Start with "Howdy!" 457 444 450 464 507 281 297 492 812 180 482
End with "Thank you" 368 280 278 260 383 177 215 170 520 131 322
Rephrase as Statement 487 584 488 388 620 232 328 496 670 127 517
Won’t Tip 485 428 413 449 454 291 338 248 672 246 480
Tip \$1 549 548 516 366 557 239 364 290 723 176 564
Tip \$10 551 559 541 368 547 239 351 286 707 178 584
Tip \$100 536 504 494 375 544 238 368 284 729 176 570
Tip \$1000 493 463 493 387 568 243 384 278 725 180 574
AIM 823 940 950 940 914 577 714 812 963 600 817
Evil Confidant 598 771 941 741 681 574 448 698 847 590 746
Refusal Suppression 631 568 788 447 651 290 544 408 743 288 812
Dev Mode v2 482 513 591 650 640 620 565 647 792 695 634
Table 14: Llama 7B number of labels changed compared to Python List per dataset for each variation.
CoLA
CoPA
ColBERT
GLUE Diagnostic
iSarcasm
IMDBSentiment
TweetStance
Jigsaw Toxicity
BoolQ
MathQA
RACE
Start with "Hello." 104 108 236 107 182 57 326 376 97 603 137
Start with "Hello!" 100 96 247 119 184 52 348 370 88 610 137
Start with "Howdy!" 107 154 322 91 236 62 319 402 96 585 111
End with "Thank you" 49 97 197 99 117 35 231 242 55 472 90
Rephrase as Statement 90 153 319 401 258 79 358 369 414 580 104
Won’t Tip 70 146 353 189 237 42 371 325 56 544 91
Tip \$1 172 115 294 123 213 47 320 312 114 648 128
Tip \$10 155 123 339 141 250 46 328 346 107 655 133
Tip \$100 146 143 454 171 334 55 383 419 105 662 136
Tip \$1000 166 154 541 235 435 64 418 462 120 666 136
AIM 525 906 786 650 642 159 930 711 376 967 440
Evil Confidant 766 970 955 930 819 477 901 633 587 967 767
Refusal Suppression 461 163 544 499 154 298 620 401 511 575 278
Dev Mode v2 162 817 491 550 312 178 581 370 445 697 546
Table 15: Llama 13B number of labels changed compared to Python List per dataset for each variation.
CoLA
ColBERT
GLUE Diagnostic
iSarcasm
IMDBSentiment
Jigsaw Toxicity
MathQA
CoPA
TweetStance
BoolQ
RACE
Start with "Hello." 33 186 118 92 23 227 168 75 267 61 128
Start with "Hello!" 51 160 117 95 30 216 182 80 219 84 143
Start with "Howdy!" 64 183 125 119 50 196 204 92 297 81 111
End with "Thank you" 181 392 85 506 66 116 127 55 113 41 79
Rephrase as Statement 66 167 375 78 36 180 211 92 346 157 125
Won’t Tip 86 258 126 77 24 183 199 84 157 74 97
Tip \$1 86 237 177 99 37 245 320 95 319 113 118
Tip \$10 115 267 145 132 38 248 353 137 371 112 118
Tip \$100 158 299 146 161 50 277 392 179 437 114 130
Tip \$1000 217 340 184 208 56 292 391 261 482 110 136
AIM 326 425 734 811 111 482 688 631 422 254 280
Evil Confidant 923 971 875 979 422 726 883 929 905 785 685
Refusal Suppression 472 803 417 602 305 402 512 202 590 282 351
Dev Mode v2 830 855 929 822 113 598 748 451 582 435 416
Table 16: Llama 70B number of labels changed compared to Python List per dataset for each variation.
@@ Appendix C No Specified Format Analysis
Figure 4: ChatGPT’s number of predictions that change (out of 11,000) compared to the No Specified Format. Red bars correspond to the number of invalid responses provided by the model.
@@
Table 17: Accuracy of each prompt variation on each task when using no specified output format on each variation. Red percentages indicate that the accuracy dropped from there baseline (No Specified Format) while green percentages indicate the accuracy increased.